### v1

@@ 2.4 Computing synthetic light curves
Synthetic light curves are calculated during the post-processing of the GCMs using prt_phasecurve222https://prt-phasecurve.readthedocs.io/en/latest/ (Schneider et al., 2022b), based on the petitRADTRANS code (Mollière et al., 2019, 2020). We first compute a detailed radiation field throughout the atmosphere, which is then interpolated to a grid of 100 points on the side of the planet facing a hypothetical observer. We then integrate the outgoing flux over those 100 points and the chosen wavelength regime, making sure to take the viewing angle into account assuming an inclination of 90°. Finally, we divide by the stellar flux in the same wavelength range. For this work, we reduced the original forty streams used in the initial radiation field computation to six streams and the wavelength resolution from 1000 to 100. This drastically reduces the computation time of the light curve while affecting the precision of a test case based on the tidally locked exoplanet HD 209458 b by less than 1%.
We generate light curves in a wavelength range of 5 – 12 m to represent the wavelength range of the Mid-InfraRed Instrument’s (MIRI) Low Resolution Spectroscopy (LRS) mode aboard JWST, which is commonly used to study exoplanets (Rieke et al., 2015; Kendrew et al., 2015), including some planets in our sample333Approved cycle 1 program #2008 A Blast From the Past: A Spectroscopic look at the Flash Heating of HD80606b. https://www.stsci.edu/jwst/science-execution/program-information?id=2008. To gain some understanding of how sensitive the light curve is to the wavelength regime, we also separate the wavelength range into four, approximately representing the F560W, F770W, F1000W and F1130W imaging filters of MIRI (Rieke et al., 2015).
@@ 3.1 Pseudo-synchronously rotating models
Each of the five eccentric exoplanets in our sample features a prominent equatorial superrotating jet throughout the entire orbit. Kataria et al. (2013); Lewis et al. (2014, 2017); Tsai et al. (2023b) found a jet on highly eccentric planets as well, demonstrating that the dynamical regime is quite robust. As in these previous works, the jet in the upper atmosphere is briefly replaced by fast day-to-night wind during the periastron passage, only to be quickly restored as the planet-star distance increases again. Lewis et al. (2017) and Tsai et al. (2023b) also note that the jet speed on HD 80606 b decreases over the orbit. Examining Fig. 2, which shows the maximal jet speed of each planet over time, we see that the jet slows down from about 1400 m s to 1000 m s on HD 80606 b, in reasonable agreement with Lewis et al. (2017); Tsai et al. (2023b). TOI-4127 b’s jet similarly slows down from 1700 m s to 1400 m s. The other three planets do not undergo such a consistent deceleration because their jets are continuously driven by the fairly strong irradiation they receive, as indicated by the equilibrium temperature shown in Fig. 3. The equilibrium temperature is the black body temperature a planet needs to radiate away the irradiation. Neglecting the albedo effect, it is
Herein, and are the effective temperature and radius of the host star and is the distance to the planet. As they are not as strongly irradiated, the thermal forcing on HD 80606 b and TOI-4127 b is insufficient to continuously feed the superrotation. Instead, their jets are only driven around the periastron passage when the irradiation is intense and the substellar point does not move as much because of the assumed pseudo-synchronous rotation period, resulting in a thermal forcing regime that strongly resembles that of a tidally locked hot Jupiter. Outside the periastron passage, the jet continues under its inertia, but is slowly decelerated by friction processes.
Figure 4 displays the contribution of the planet to the total brightness of each system in different wavelength ranges. The solid black line shows the integral of the full 5 – 12 m range. The planet’s relative contribution is greater at longer wavelengths because of its cooler temperature, which is especially notable on HD 80606 b, TOI-4127 b and HD 17156 b. Each planet’s global maximum in emission occurs shortly after the periastron passage, except for HD 80606 b’s because of its illuminated side being in the observer’s view before the periastron passage, as shown in Fig. 1. After this initial periastron peak, the brightness of HAT-P-2 b and TOI-3362 b decreases mostly monotonically until the next periastron passage. TOI-4127 b, on the other hand, shows one local maximum after the first peak while HD 80606 b has four. These ringing peaks are especially high at longer wavelengths. HD 17156 b has local maxima with small amplitudes only at longer wavelengths.
@@ 3.2 Breaking pseudo-synchronicity on HD 80606 b
To demonstrate the influence of the planet’s rotation period, we repeat the HD 80606 b model with a rotation period equal to half and double the pseudo-synchronous period. Figure 5 shows how the speed of the equatorial jet differs from the pseudo-synchronous case. We note that a longer rotation period leads to a significantly faster equatorial jet. This is explained by the substellar point moving eastwards during the periastron passage if the rotation period is longer than pseudo-synchronous. This results in more irradiation getting deposited into the already hot part of the atmosphere, thereby increasing the temperature gradient and thus the geostrophic wind speed, resulting in stronger angular momentum transport to the equator and hence a faster jet (Showman et al., 2009, 2015). Moreover, the jet is broader with a longer rotation period as explained by the equatorial Rossby deformation length , a measure of the typical jet width given by
with the Brunt-Väisälä frequency, Boltzmann’s constant, the temperature, the mean molecular weight, the atomic mass and the surface gravity (Showman & Guillot, 2002). Consequently, the re-distribution of heat becomes far more efficient with a longer rotation period. When decreasing the rotation period, the opposite occurs. The substellar point rapidly moves west at all times, even during the periastron passage, so the irradiation is spread thin and the resulting weak temperature gradient fails to pump sufficient angular momentum to the equator.
Fig. 6 demonstrates the importance of the planetary rotation period to the ringing phenomenon, as the case with a longer rotation period displays only one strong peak for two reasons. Firstly, heat is distributed more effectively by the circulation, which reduces the temperature contrast between the flash-heated and unheated hemispheres and hence the ringing amplitude. Secondly, as the rotation period is slower, it takes longer for the heated side to come into view, allowing it more time to cool and for the heat to be distributed. For the opposite reasons, the case with a short rotation period shows ten local maxima, the first of which are taller than in the pseudo-synchronously rotating case. Interestingly, the rapidly rotating case still shows small peaks over 8 d after the periastron passage, whereas the pseudo-synchronous case’s last local maximum occurs about 6 d after the periastron passage, demonstrating that the lowered efficiency of heat redistribution by circulation plays an important role.
@@ 4 Discussion
4.1 Timing of ringing peaks
Due to the strong increase in heating, the emission of the planet peaks sharply around the periastron passage on highly eccentric exoplanets. The timing of this peak relative to the periastron passage is affected by a number of parameters, including the planet’s rotation period, thereby offering a rare opportunity for an observational constraint on planetary rotation, a route taken by de Wit et al. (2016) and Lewis et al. (2017). Consider as an example the emission of HAT-P-2 b, where about half of the dayside is in the viewing disk of an observer on Earth, as visualised in Fig. 1. In that case, one might expect the dayside to come entirely into view after h. However, our synthetic light curve shown in Fig. 4 predicts that the maximum in observed emission would occur h after the periastron passage. Here we ought to include the effect of the superrotating equatorial jet which pushes the gas heated during the heating flash eastwards, hence bringing it into view earlier. We may roughly estimate the jet-enhanced timing of the periastron peak to be
where is the eastward speed at the equator due to rotation and the jet speed after periastron. For HAT-P-2 b, we estimate m s from Fig. 2 and m s, resulting in h. This matches the predicted peak timing of h given the temporal resolution of about h.
@@ 4.1 Timing of ringing peaks
with the heat capacity at constant pressure and the Stefan-Boltzmann constant (Showman & Guillot, 2002). On TOI-4127 b for instance, we estimate the maximum to occur h after the periastron passage using Eq. (4), while the synthetic light curve predicts h after the periastron passage. As such, the cooling after periastron passage and the degeneracy between rotation and the superrotating jet seriously complicate the extraction of the rotation period from the periastron peak, leading to large uncertainties in studies such as de Wit et al. (2016).
Under the right circumstances, ringing can isolate the rotation period from the timing of the ringing peaks. In the HD 80606 b simulation, the spacings between the peaks are d, d, d and d with a resolution of about d while the rotation period is d. The spacing between the periastron peak and first and second ringing peaks is shorter than the rotation period. To account for the superrotating jet, we use
Using m s and m s for HD 80606 b, we estimate d. This underestimates the modelled peak spacings in part because of the thinness of the jet due to HD 80606 b’s short rotation period and low temperature reducing the equatorial Rossby deformation radius (see Eq. (3)). Figure 7 shows the temperature evolution on HD 80606 b after periastron passage at 1 bar as seen by an observer on Earth. Initially, the flash heated gas is concentrated in one hemisphere. As the thin equatorial jet drags the hot gas at the equator eastwards, the spacing between ringing peaks is smaller than the rotation period. However, as the jet is quite thin, some flash-heated gas remains to the north-east and south-east of the substellar point at the periastron, indicated by the green dot. After an advection timescale given by
, the temperature on the equator is mostly homogenised by the jet. On HD 80606 b we estimate d, matching the fourth panel of Fig. 7. Yet the parcels of flash-heated gas outside the jet still maintain their higher temperature. As that gas continues to rotate in and out of view the last two ringing peaks follow with a spacing consistent with the rotation period. Consequently, the degeneracy between the rotation period and wind speed is broken, so the rotation period can be directly measured from the spacing between the last peaks.
@@ 4.2 Detecting ringing on these five exoplanets
Lewis et al. (2017) demonstrated that clouds may seriously diminish ringing on HD 80606 b. Since our model does not include the effects of clouds, we likely overestimate the amplitude of the ringing peaks. Similarly, Tsai et al. (2023b) show that including photo-chemistry in the post-processing of their HD~80606~b models slightly reduces the emission magnitude after the periastron passage. However, unlike clouds this is unlikely to qualitatively alter the light curve as the composition at photospheric altitudes returns close to pre-periastron levels by the time the first ringing peak occurs. On top of the lack of clouds, our model also assumes an obliquity of 0. HD 80606 b, TOI-4127 b and HD 17156 b have an orbital period greater than the 20 d limit beyond which one cannot safely assume tides to have aligned the rotational and orbital axes (Rauscher et al., 2023). Using a simple 2D shallow-water model, Ohno & Zhang (2019a, b) showed how the obliquity and corresponding viewing geometry can strongly affect the light curve, which was confirmed by Rauscher et al. (2023). If, for instance, one of the poles receives most of the flash-heating at the periastron, that heat would not be brought in and out of view by rotation, hence no ringing. This too points towards our synthetic light curves overestimating the amplitude of the ringing. If little or no ringing is detected on HD 80606 b, this could be explained by a rotation period longer than pseudo-synchronous, clouds or obliquity. If clouds were responsible, one might expect to detect them using transit spectroscopy. If the non-detection of ringing were due to a long rotation period or an oblique orbit, the planet would prove itself useful to the research of tidal evolution and possibly to the detection of obliquity using secondary eclipse mapping (Rauscher, 2017; Rauscher et al., 2023).
@@
In our synthetic light curve of HD 80606 b, the final ringing peak takes place 6.5 d after the periastron passage. To include the decline after that local maximum, one would need a baseline stretching until approximately 7 d after the periastron passage. As the global maximum occurs shortly before the periastron passage, the baseline should be approximately 8 d long to cover all peaks. This baseline is considerably longer than those of currently accepted proposals for JWST observations of HD 80606 b which were allotted a prime time of approximately one day444Approved cycle 1 program #2008 A Blast From the Past: A Spectroscopic look at the Flash Heating of HD80606b. https://www.stsci.edu/jwst/science-execution/program-information?id=2008 and Approved cycle 1 program #2488 Real Time Exoplanet Meteorology: Direct Measurement of Cloud Dynamics on the High-Eccentricity Hot Jupiter HD80606 b. https://www.stsci.edu/jwst/science-execution/program-information?id=2488. As the amplitude of the ringing in our synthetic light curves is likely overestimated, it is uncertain whether the fourth ringing peak is indeed detectable. Therefore, it may be reasonable to shorten the baseline to omit the fourth ringing peak, in which case the baseline would stretch until approximately 5 d after the periastron passage. Alternatively, a weaker constraint can be placed on the rotation period through a comparison of observations to GCMs with different rotation periods as in Fig. 6. To clearly distinguish the case from the case of HD 80606 b, a run ought to last for about 3 d after the periastron when the observed emission on the case increases while the emission continues to decrease.
@@ 4.3 Why only HD 80606 b displays ringing
In our sample of five highly eccentric exoplanets, no planet displays any detectable ringing peaks, except for HD 80606 b which has four. This remarkable difference in the level of ringing may be explained by a number of factors. Firstly, the extremely eccentric orbit of HD 80606 b results in a great difference in irradiation at the periastron compared to throughout most of the orbit, resulting in a large temperature contrast. As ringing is the consequence of a difference in emission between two hemispheres, a high eccentricity increases the amplitude of ringing peaks. Furthermore, under the assumption that the planet rotates pseudo-synchronously, a high eccentricity leads to a short rotation period, meaning little of the heat is distributed by circulation or irradiated away between rotations. Indeed, HD 80606 b has the second shortest pseudo-synchronous rotation period in our sample (see Table 1). However, Kataria et al. (2013) found an increase in the speed of the equatorial jet with eccentricity, which can raise the effectiveness of the circulation. Nevertheless, their work still suggests that a high eccentricity is beneficial for ringing, as they found no ringing on a hypothetical planet with and strong ringing on a similar case with .
Secondly, ringing in our sample is anti-correlated with the mean equilibrium temperature shown in Fig. 3. HD 80606 b has the lowest mean equilibrium temperature, while the two most irradiated cases, HAT-P-2 b and TOI-3362 b do not display any ringing. This is in part due to the long radiative cooling timescale of cool gas, which helps to maintain the temperature contrast between the two hemispheres. Furthermore, weaker irradiation is known to lead to a slower equatorial jet and hence less efficient heat redistribution (Kataria et al., 2013; Showman et al., 2015). However, Kataria et al. (2013) noted that a shorter orbital period and hence a greater equilibrium temperature leads to more and higher ringing peaks in their model with . They indicate that the strengthened ringing is aided by the shortened pseudo-synchronous rotation period. As such, the net effect of the orbital period and mean equilibrium temperature is still unclear.

### v2

@@ 2.4 Computing synthetic light curves
Synthetic light curves were calculated during the post-processing of our GCMs using prt_phasecurve222https://prt-phasecurve.readthedocs.io/en/latest/ (Schneider et al., 2022b), based on the petitRADTRANS code (Mollière et al., 2019, 2020). We first computed a detailed radiation field throughout the atmosphere, which was then interpolated to a grid of 100 points on the side of the planet facing a hypothetical observer. We then integrated the outgoing flux over those 100 points and the chosen wavelength regime, making sure to take the viewing angle into account assuming an inclination of 90°. Finally, we divided by the stellar flux in the same wavelength range. For this work, we reduced the original forty streams used in the initial radiation field computation to six streams and the wavelength resolution from 1000 to 100. This drastically reduces the computation time of the light curve while affecting the precision of a test case based on the tidally locked exoplanet HD 209458 b by less than 1%.
We generated light curves in a wavelength range of 5 – 12 m to represent the wavelength range of the Mid-InfraRed Instrument (MIRI) low resolution spectroscopy (LRS) mode aboard JWST. This instrument is commonly used to study exoplanets (Rieke et al., 2015; Kendrew et al., 2015), including some planets in our sample333Approved cycle 1 programme #2008 A Blast From the Past: A Spectroscopic look at the Flash Heating of HD80606b. https://www.stsci.edu/jwst/science-execution/program-information?id=2008. To gain some understanding of how sensitive the light curve is to the wavelength regime, we also separated the wavelength range into four sub-ranges, which approximately represent the F560W, F770W, F1000W, and F1130W imaging filters of MIRI (Rieke et al., 2015).
@@ 3.1 Pseudo-synchronously rotating models
Each of the five eccentric exoplanets in our sample features a prominent equatorial super-rotating jet throughout the entire orbit. Kataria et al. (2013), Lewis et al. (2014, 2017), and Tsai et al. (2023b) find a jet on highly eccentric planets as well, demonstrating that the dynamical regime is quite robust. As in these previous works, the jet in the upper atmosphere is briefly replaced by fast day-to-night wind during the periastron passage, only to be quickly restored as the planet-star distance increases again. Lewis et al. (2017) and Tsai et al. (2023b) also note that the jet speed on HD 80606 b decreases over the orbit. Examining Fig. 2, which shows the maximal jet speed of each planet over time, we see that the jet slows down from about 1400 m s-1 to 1000 m s-1 on HD 80606 b, in reasonable agreement with Lewis et al. (2017) and Tsai et al. (2023b). TOI-4127 b’s jet similarly slows down from 1700 m s-1 to 1400 m s-1. The other three planets do not undergo such a consistent deceleration because their jets are continuously driven by the fairly strong irradiation they receive, as indicated by the equilibrium temperature shown in Fig. 3. The equilibrium temperature is the black body temperature a planet needs to radiate away the irradiation. Neglecting the albedo effect, it is
Herein, and are the effective temperature and radius of the host star and is the distance to the planet. As they are not as strongly irradiated, the thermal forcing on HD 80606 b and TOI-4127 b is insufficient to continuously feed the super-rotation. Instead, their jets are only driven around the periastron passage when the irradiation is intense and the substellar point does not move as much because of the assumed pseudo-synchronous rotation period, resulting in a thermal forcing regime that strongly resembles that of a tidally locked hot Jupiter. Outside the periastron passage, the jet continues under its inertia, but is slowly decelerated by friction processes.
Figure 4 displays the contribution of the planet to the total brightness of each system in different wavelength ranges. The solid black line shows the integral of the full 5 – 12 m range. The planet’s relative contribution is greater at longer wavelengths because of its cooler temperature, which is especially notable on HD 80606 b, TOI-4127 b, and HD 17156 b. Each planet’s global maximum in emission occurs shortly after the periastron passage, except for HD 80606 b’s due to its illuminated side being in the observer’s view before the periastron passage, as shown in Fig. 1. After this initial periastron peak, the brightness of HAT-P-2 b and TOI-3362 b decreases mostly monotonically until the next periastron passage. TOI-4127 b, on the other hand, shows one local maximum after the first peak while HD 80606 b has four. These ringing peaks are especially high at longer wavelengths. HD 17156 b has local maxima with small amplitudes only at longer wavelengths.
@@ 3.2 Breaking pseudo-synchronicity on HD 80606 b
To demonstrate the influence of the planet’s rotation period, we repeated the HD 80606 b model with a rotation period equal to half and double the pseudo-synchronous period. Figure 5 shows how the speed of the equatorial jet differs from the pseudo-synchronous case. We note that a longer rotation period leads to a significantly faster equatorial jet. This is explained by the substellar point moving eastwards during the periastron passage if the rotation period is longer than pseudo-synchronous. This leads to more irradiation getting deposited into the already hot part of the atmosphere, thereby increasing the temperature gradient and thus the geostrophic wind speed, resulting in stronger angular momentum transport to the equator and hence a faster jet (Showman et al., 2009, 2015). Moreover, the jet is broader with a longer rotation period as explained by the equatorial Rossby deformation length , a measure of the typical jet width given by
with the Brunt-Väisälä frequency, Boltzmann’s constant, the temperature, the mean molecular weight, the atomic mass, and the surface gravity (Showman & Guillot, 2002). Consequently, the redistribution of heat becomes far more efficient with a longer rotation period. When decreasing the rotation period, the opposite occurs. The substellar point rapidly moves west at all times, even during the periastron passage, so the irradiation is spread thin and the resulting weak temperature gradient fails to pump sufficient angular momentum to the equator.
Figure 6 demonstrates the importance of the planetary rotation period to the ringing phenomenon, as the case with a longer rotation period displays only one strong peak for two reasons. Firstly, heat is distributed more effectively by the circulation, which reduces the temperature contrast between the flash-heated and unheated hemispheres and hence the ringing amplitude. Secondly, as the rotation period is longer, it takes longer for the heated side to come into view, allowing it more time to cool and for the heat to be distributed. For the opposite reasons, the case with a short rotation period shows ten local maxima, the first of which are taller than in the pseudo-synchronously rotating case. Interestingly, the rapidly rotating case still shows small peaks over 8 d after the periastron passage, whereas the pseudo-synchronous case’s last local maximum occurs about 6 d after the periastron passage, demonstrating that the lowered efficiency of heat redistribution by circulation plays an important role.
@@ 4 Discussion
4.1 Timing of the ringing peaks
Due to the strong increase in heating, the emission of the planet peaks sharply around the periastron passage on highly eccentric exoplanets. The timing of this peak relative to the periastron passage is affected by a number of parameters, including the planet’s rotation period, thereby offering a rare opportunity for an observational constraint on planetary rotation, a route taken by de Wit et al. (2016) and Lewis et al. (2017). Consider as an example the emission of HAT-P-2 b, where about half of the dayside is in the viewing disk of an observer on Earth, as visualised in Fig. 1. In that case, one might expect the dayside to come entirely into view after h. However, our synthetic light curve shown in Fig. 4 predicts that the maximum in observed emission would occur h after the periastron passage. Here we ought to include the effect of the super-rotating equatorial jet, which pushes the gas heated during the heating flash eastwards, hence bringing it into view earlier. We can roughly estimate the jet-enhanced timing of the periastron peak to be
where is the eastward speed at the equator due to rotation and the jet speed after periastron. For HAT-P-2 b, we estimate m s-1 from Fig. 2 and m s-1, resulting in h. This matches the predicted peak timing of h given the temporal resolution of about h.
@@ 4.1 Timing of the ringing peaks
with the heat capacity at constant pressure and the Stefan-Boltzmann constant (Showman & Guillot, 2002). On TOI-4127 b for instance, we estimate the maximum to occur h after the periastron passage using Eq. (4), while the synthetic light curve predicts h after the periastron passage. As such, the cooling after periastron passage and the degeneracy between rotation and the super-rotating jet seriously complicate the extraction of the rotation period from the periastron peak, leading to large uncertainties in studies such as de Wit et al. (2016).
Under the right circumstances, ringing can isolate the rotation period from the timing of the ringing peaks. In the HD 80606 b simulation, the spacings between the peaks are d, d, d, and d with a resolution of about d, and the rotation period is d. The spacing between the periastron peak and first and second ringing peaks is shorter than the rotation period. To account for the super-rotating jet, we used
Using m s-1 and m s-1 for HD 80606 b, we estimate d. This underestimates the modelled peak spacings in part because of the thinness of the jet due to HD 80606 b’s short rotation period and low temperature reducing the equatorial Rossby deformation radius (see Eq. (3)). Figure 7 shows the temperature evolution on HD 80606 b after periastron passage at 1 bar as seen by an observer on Earth. Initially, the flash heated gas is concentrated in one hemisphere. As the thin equatorial jet drags the hot gas at the equator eastwards, the spacing between ringing peaks is smaller than the rotation period. However, as the jet is quite thin, some flash-heated gas remains to the north-east and south-east of the substellar point at the periastron, indicated by the green dot. After an advection timescale given by
the temperature on the equator is mostly homogenised by the jet. On HD 80606 b we estimate d, matching the fourth panel of Fig. 7. Yet the parcels of flash-heated gas outside the jet still maintain their higher temperature. As that gas continues to rotate in and out of view the last two ringing peaks follow with a spacing consistent with the rotation period. Consequently, the degeneracy between the rotation period and wind speed is broken, so the rotation period can be directly measured from the spacing between the last peaks.
@@ 4.2 Detecting ringing on these five exoplanets
Lewis et al. (2017) demonstrate that clouds may seriously diminish ringing on HD 80606 b. Since our model does not include the effects of clouds, we likely overestimate the amplitude of the ringing peaks. Similarly, Tsai et al. (2023b) show that including photo-chemistry in the post-processing of their HD~80606~b models slightly reduces the emission magnitude after the periastron passage. However, unlike clouds this is unlikely to qualitatively alter the light curve as the composition at photospheric altitudes returns close to pre-periastron levels by the time the first ringing peak occurs. On top of the lack of clouds, our model also assumes an obliquity of 0. HD 80606 b, TOI-4127 b, and HD 17156 b have an orbital period greater than the 20 d limit beyond which one cannot safely assume tides to have aligned the rotational and orbital axes (Rauscher et al., 2023). Using a simple 2D shallow-water model, Ohno & Zhang (2019a, b) show how the obliquity and corresponding viewing geometry can strongly affect the light curve, which is confirmed by Rauscher et al. (2023). If, for instance, one of the poles receives most of the flash-heating at the periastron, that heat would not be brought in and out of view by rotation, and hence no ringing. This too points towards our synthetic light curves overestimating the amplitude of the ringing. If little or no ringing is detected on HD 80606 b, this could be explained by a rotation period longer than pseudo-synchronous, clouds or obliquity. If clouds were responsible, one might expect to detect them using transit spectroscopy. If the non-detection of ringing were due to a long rotation period or an oblique orbit, the planet would prove itself useful to the research of tidal evolution and possibly to the detection of obliquity using secondary eclipse mapping (Rauscher, 2017; Rauscher et al., 2023).
@@
In our synthetic light curve of HD 80606 b, the final ringing peak takes place 6.5 d after the periastron passage. To include the decline after that local maximum, one would need a baseline stretching until approximately 7 d after the periastron passage. As the global maximum occurs shortly before the periastron passage, the baseline should be approximately 8 d long to cover all peaks. This baseline is considerably longer than those of currently accepted proposals for JWST observations of HD 80606 b, which were allotted a prime time of approximately one day444Approved cycle 1 programme #2008, A Blast From the Past: A Spectroscopic look at the Flash Heating of HD80606b. https://www.stsci.edu/jwst/science-execution/program-information?id=2008 and Approved cycle 1 program #2488, Real Time Exoplanet Meteorology: Direct Measurement of Cloud Dynamics on the High-Eccentricity Hot Jupiter HD80606 b. https://www.stsci.edu/jwst/science-execution/program-information?id=2488. As the amplitude of the ringing in our synthetic light curves is likely overestimated, it is uncertain whether the fourth ringing peak is indeed detectable. Therefore, it may be reasonable to shorten the baseline to omit the fourth ringing peak, in which case the baseline would stretch until approximately 5 d after the periastron passage. Alternatively, a weaker constraint can be placed on the rotation period through a comparison of observations to GCMs with different rotation periods as in Fig. 6. To clearly distinguish the case from the case of HD 80606 b, a run ought to last for about 3 d after the periastron when the observed emission on the case increases while the emission continues to decrease.
@@ 4.3 Why only HD 80606 b displays ringing
In our sample of five highly eccentric exoplanets, no planet displays any detectable ringing peaks except HD 80606 b, which has four. This remarkable difference in the level of ringing may be explained by a number of factors. Firstly, the extremely eccentric orbit of HD 80606 b results in a great difference in irradiation at the periastron compared to throughout most of the orbit, resulting in a large temperature contrast. As ringing is the consequence of a difference in emission between two hemispheres, a high eccentricity increases the amplitude of ringing peaks. Furthermore, under the assumption that the planet rotates pseudo-synchronously, a high eccentricity leads to a short rotation period, meaning little of the heat is distributed by circulation or irradiated away between rotations. Indeed, HD 80606 b has the second shortest pseudo-synchronous rotation period in our sample (see Table 1). However, Kataria et al. (2013) find an increase in the speed of the equatorial jet with eccentricity, which can raise the effectiveness of the circulation. Nevertheless, their work still suggests that a high eccentricity is beneficial for ringing, as they found no ringing on a hypothetical planet with and strong ringing on a similar case with .
Secondly, ringing in our sample is anti-correlated with the mean equilibrium temperature shown in Fig. 3. HD 80606 b has the lowest mean equilibrium temperature, while the two most irradiated cases, HAT-P-2 b and TOI-3362 b do not display any ringing. This is in part due to the long radiative cooling timescale of cool gas, which helps maintain the temperature contrast between the two hemispheres. Furthermore, weaker irradiation is known to lead to a slower equatorial jet and hence less efficient heat redistribution (Kataria et al., 2013; Showman et al., 2015). However, Kataria et al. (2013) note that a shorter orbital period and hence a greater equilibrium temperature leads to more and higher ringing peaks in their model with . They indicate that the strengthened ringing is aided by the shortened pseudo-synchronous rotation period. As such, the net effect of the orbital period and mean equilibrium temperature is still unclear.